OpenStudy (anonymous):

Divide and simplify:d/b^2/d^2/b^3

OpenStudy (anonymous):

d over B^2 divided by d^2over b^3

OpenStudy (anonymous):

like this? \[\frac{\frac{d}{b^{2}}}{\frac{d^{2}}{b^{3}}}\] looks funny lol

OpenStudy (anonymous):

I am having a hard time with this stuff.Math is not my strong point

myininaya (myininaya):

\[\frac{d}{b^2} \div \frac{d^2}{b^3}=\frac{d}{b^2}*\frac{b^3}{d^2}\]

OpenStudy (anonymous):

yeah i should have written it like that >.<

myininaya (myininaya):

\[\frac{db^3}{b^2d^2}=\frac{d}{d^2}*\frac{b^3}{b^2}\]

myininaya (myininaya):

do you see anything that cancels?

myininaya (myininaya):

\[\frac{d}{d*d}*\frac{b^2*b}{b^2}\]

OpenStudy (anonymous):

says its wrong

OpenStudy (anonymous):

never mind

myininaya (myininaya):

we haven't simplified do you anything that cancels?

OpenStudy (anonymous):

so the d's cancel out?

myininaya (myininaya):

a pair of d's yes anything else?

myininaya (myininaya):

\[\frac{1}{d}*\frac{b^2*b}{b^2}\]

OpenStudy (anonymous):

okay and b^2

myininaya (myininaya):

yes \[\frac{1}{d}*b=\frac{b}{d}\]

OpenStudy (anonymous):

So the answer is going to be B over D

myininaya (myininaya):

b/d

OpenStudy (anonymous):

thank you